# Optimus: Accelerating Large-Scale Multi-Modal LLM Training by Bubble Exploitation

### v1

@@ 5.2.3 Multi-Encoder MLLM Experiment
Table 6: Multi-encoder MLLM configurations
@@
Figure 16: Training performance of Optimus and Megatron-LM on multi-encoder MLLMs
@@ 5.3.1 Optimus Memory
Figure 17: GPU memory usage of Optimus and Megatron-based baselines for MLLMs shown in Table 3
@@ 5.3.2 Bubble Scheduler Algorithm
Table 7: Scheduling efficiency and algorithm runtime of the bubble scheduler algorithm
@@ 6 Discussion
Complex computation graph. Optimus focuses on the bubble scheduling on typical MLLM model architecture, which consists of multimodal encoders followed by one LLM. We may further explore the bubble scheduling for complex MLLM computation graphs. A new partitioning algorithm is required to divide the computation graph into the backbone pipeline schedule and the bubble-filling workload. And the bubble scheduling algorithm of Optimus can be easily extended to the partitioned computation graph.
Other pipeline schedules. We use a widely-used Megatron-LM interleaved 1F1B pipeline schedule for MLLM training. However, there exist other pipeline schedules (e.g., Chimera [17] and zero-bubble pipeline [24]) that may have superior performance in certain scenarios. The bubble scheduling of Optimus is orthogonal to these pipeline schedule optimizations, and Optimus can be applied to other pipeline schedules when the specific encoder-LLM dependency is analyzed and addressed.
Online scheduling. For simplicity, our bubble scheduling algorithm omits the consideration of fluctuating runtime execution time of CUDA kernels. We collect performance statistics such as CUDA kernel execution time to identify the bubble occurrence and duration during a training step for the bubble scheduler, assuming the behavior remains the same in the following training steps. The bubble scheduling may be sub-optimal when there is a significant deviation in the predicted pipeline execution time. For instance, the insertion of encoder computation into a non-bubble position may result in larger, unexpected pipeline bubbles with altered execution orders. A potential solution is to use real-time performance monitoring, and dynamically fine-tune the bubble scheduling.
@@ 7 Related works
Multi-modal training. Pytorch FSDP training [37] supports only data parallelism and is less efficient than hybrid parallel strategies. Alpa [38] automates parallelism for various models but falls short by not supporting state-of-the-art 1F1B-interleave pipeline parallelism [21] and requiring more memory than the optimized Megatron-LM framework [26], also missing opportunities in pipeline optimization due to its unified view of encoders and decoders. DistMM [13] provides solutions to orchestrating multiple parallel encoders but it is designed for contrastive learning and overlooks the decoder, leaving a gap in comprehensive training efficiency.
Bubble reducing. Previous efforts in reducing “bubbles” have approached the problem from various angles. The 1F1B-interleave pipeline [21] technique minimizes bubbles by chunking the model and alternating these chunks across different stages, whereas the Zero bubble pipeline [24] approach further granulates backward pass computations to eliminate bubbles. However, in practice, the Zero bubble pipeline schedule cannot completely remove all pipeline bubbles because it requires changes to the optimizer, which raises concerns about end-to-end model convergence. Sarathi [3] splits sequence into smaller chunks to do multi-step prefilling and thus reduce pipeline bubble in LLM inference. On the other hand, asynchronous tensor parallelism [27] and Google’s overlapping technique [32] aim to overlap tensor parallelism communication with computation but are limited by specific hardware configurations and struggle to maintain full overlap as computing capabilities advance.
Bubble exploiting. Pipefisher [23] leverages pipeline bubbles across multiple training steps to complete the K-FAC, whereas our method operates within a single synchronized training step, focusing on immediate optimization. Hydro’s Bubble Squeezer [12] utilizes GPT model bubbles for independent tasks like hyperparameter tuning which can not enhance the performance of the training steps themselves. Bamboo [30] employs pipeline bubbles for redundant computations to mitigate the impact of preemption in training on volatile instances, based on the assumption that later pipeline stages host more layers, which often does not hold in large language model (LLM) training scenarios.
@@ 8 Conclusion
We present Optimus, a distributed MLLM training system that enables the scheduling of encoder computation within LLM bubbles to reduce end-to-end MLLM training time. To reduce GPU bubbles during MLLM training, Optimus partitions multimodal encoders and the LLM backbone, and schedules encoder computation in LLM bubbles. We search for the optimal parallelism plan for the encoders with the consideration of memory and computation resource constraints, which balances the encoder computation among GPUs for bubble filling. Optimus further employs a bubble scheduling algorithm to address encoder-LLM dependency and select the optimal schedule for filling kernel-level encoder computation into sub-millisecond LLM bubbles. Our extensive experiments demonstrate that Optimus can accelerate MLLM training by 20.5%-21.3% with ViT-22B and GPT-175B model over 3072 GPUs compared to baselines and significantly outperforms existing MLLM training systems by 20.3% on average.

### v2

@@ 5.2.3 Multi-Encoder MLLM Experiment
Table 6: Multi-encoder MLLM configurations.
@@
Figure 16: Training performance of Optimus and Megatron-LM on multi-encoder MLLMs.
@@ 5.3.1 Optimus Memory
Figure 17: GPU memory usage of Optimus and Megatron-based baselines for MLLMs shown in Table 3.
@@ 5.3.2 Bubble Scheduler Algorithm
Table 7: Scheduling efficiency and algorithm runtime of the bubble scheduler algorithm.
@@ 6 Discussion
MLLM training with frozen parameters. Optimus readily supports multi-stage training workflows commonly used in practice, such as those adopted by LLaVA [18]. While our paper focuses on the general setting where all parameters—including those of the encoder and LLM—are updated, Optimus can naturally adapt to stages where only adapters are trained. In such cases, Optimus schedules the encoder + adapter forward pass and the adapter’s backward pass into the LLM pipeline bubbles, while skipping the encoder’s backward computation due to frozen parameters. This maintains correct data dependencies and continues to utilize bubbles effectively.
Complex computation graph. Optimus focuses on the bubble scheduling on a typical MLLM model architecture, which consists of multimodal encoders followed by one LLM. We may further explore the bubble scheduling for complex MLLM computation graphs. A new partitioning algorithm is required to divide the computation graph into the backbone pipeline schedule and the bubble-filling workload. And the bubble scheduling algorithm of Optimus can be easily extended to the partitioned computation graph.
Other pipeline schedules. We use a widely used Megatron-LM interleaved 1F1B pipeline schedule for MLLM training. However, there exist other pipeline schedules (e.g., Chimera [16] and zero-bubble pipeline [23]) that may have superior performance in certain scenarios. The bubble scheduling of Optimus is orthogonal to these pipeline schedule optimizations, and Optimus can be applied to other pipeline schedules when the specific encoder-LLM dependency is analyzed and addressed.
Online scheduling. Our bubble scheduling algorithm simplifies execution by omitting the consideration of fluctuations in CUDA kernel runtime. We collect performance statistics, such as kernel execution time, to detect bubbles during a training step, assuming consistent behavior in future steps. However, deviations from predicted execution times can lead to suboptimal scheduling, with larger or altered pipeline bubbles. A possible solution is real-time performance monitoring to dynamically adjust the schedule.
Exclusion of DiffusionPipe and DistTrain as Baselines. We explicitly exclude DiffusionPipe[30] and DistTrain [37] from our set of baselines for the following reasons. DiffusionPipe is specifically designed for diffusion models and has been evaluated only on small-scale clusters ($\leq$64 GPUs). Its focus on a different model family and scale makes it ill-suited for comparison with Optimus, which targets large-scale multimodal LLM (MLLM) training across thousands of GPUs. DistTrain relies on a simplified model partitioning strategy that can result in significant pipeline imbalance when applied to the heterogeneous structures of MLLMs. Furthermore, DistTrain is not open source, preventing direct empirical comparisons.
@@ 7 Related works
Multi-modal training. Pytorch FSDP training [38] supports only data parallelism and is less efficient than hybrid parallel strategies. Alpa [39] automates parallelism for various models but falls short by not supporting state-of-the-art 1F1B-interleave pipeline parallelism [20] and requiring more memory than the optimized Megatron-LM framework [25], also missing opportunities in pipeline optimization due to its unified view of encoders and decoders. DistMM [12] provides solutions to orchestrating multiple parallel encoders, but it is designed for contrastive learning and overlooks the decoder, leaving a gap in comprehensive training efficiency. DiffusionPipe [30] and DistTrain [37] are two additional works on multi-modal training, each with limitations outlined in the previous section.
Bubble reducing. Previous efforts in reducing “bubbles” have approached the problem from various angles. The 1F1B-interleave pipeline [20] technique minimizes bubbles by chunking the model and alternating these chunks across different stages, whereas the Zero bubble pipeline [23] approach further granulates backward pass computations to eliminate bubbles. However, in practice, the Zero bubble pipeline schedule cannot completely remove all pipeline bubbles because it requires changes to the optimizer, which raises concerns about end-to-end model convergence. On the other hand, asynchronous tensor parallelism [26] and Google’s overlapping technique [32] aim to overlap tensor parallelism communication with computation, but are limited by specific hardware configurations and struggle to maintain full overlap as computing capabilities advance.
Bubble exploiting. Pipefisher [22] leverages pipeline bubbles across multiple training steps to complete the K-FAC, whereas our method operates within a single synchronized training step, focusing on immediate optimization. Hydro’s Bubble Squeezer [11] utilizes GPT model bubbles for independent tasks like hyperparameter tuning, which can not enhance the performance of the training steps themselves. Bamboo [29] employs pipeline bubbles for redundant computations to mitigate the impact of preemption in training on volatile instances, based on the assumption that later pipeline stages host more layers, which often does not hold in large language model (LLM) training scenarios.
@@ 8 Conclusion
We present Optimus, a distributed MLLM training system that enables the scheduling of encoder computation within LLM bubbles to reduce end-to-end MLLM training time. To reduce GPU bubbles during MLLM training, Optimus partitions multimodal encoders and the LLM backbone, and schedules encoder computation in LLM bubbles. We search for the optimal parallelism plan for the encoders with the consideration of memory and computation resource constraints, which balances the encoder computation among GPUs for bubble filling. Optimus further employs a bubble scheduling algorithm to address encoder-LLM dependency and select the optimal schedule for filling kernel-level encoder computation into sub-millisecond LLM bubbles. Our extensive experiments demonstrate that Optimus can accelerate MLLM training by 20.5%-21.3% with ViT-22B and GPT-175B models over 3072 GPUs compared to baselines and significantly outperforms existing MLLM training systems by 20.3% on average.